# Attention-Guided Multi-scale Interaction Network for Face Super-Resolution

### v2

@@ Abstract
Recently, CNN and Transformer hybrid networks demonstrated excellent performance in face super-resolution (FSR) tasks. Since numerous features at different scales in hybrid networks, how to fuse these multi-scale features and promote their complementarity is crucial for enhancing FSR. However, existing hybrid network-based FSR methods ignore this, only simply combining the Transformer and CNN. To address this issue, we propose an attention-guided Multi-scale interaction network (AMINet), which contains local and global feature interactions and encoder-decoder phase feature interactions. Specifically, we propose a Local and Global Feature Interaction Module (LGFI) to promote fusions of global features and different receptive fields’ local features extracted by our Residual Depth Feature Extraction Module (RDFE). Additionally, we propose a Selective Kernel Attention Fusion Module (SKAF) to adaptively select fusions of different features within LGFI and encoder-decoder phases. Our above design allows the free flow of multi-scale features from within modules and between encoder and decoder, which can promote the complementarity of different scale features to enhance FSR. Comprehensive experiments confirm that our method consistently performs well with less computational consumption and faster inference.
@@
However, while existing hybrid-based FSR methods consider utilizing features from different scales to improve FSR, they ignore the problem of how we fuse these multi-scale features to make their properties better complement each other. For example, Faceformer [8] simply parallelizes the connected CNN modules and the window-based Transformer [9] modules. SCTANet [10] also only juxtaposes spatial attention-based residual blocks and multi-head self-attention in designed modules. CTCNet [5] simply connects the CNN module in tandem with the Transformer module. These methods overlook the importance of blending multi-scale features in a complementary manner and enabling smooth information flow across different scales to refine facial details effectively.
To address this problem, we propose an Attention-Guided Multi-scale Interaction Network (AMINet) for FSR in this work. Our AMINet fuses multi-scale features in two main ways, including the fusion of features obtained from self-attention and convolution and the fusion of features at different stages of the encoder-decoder. Specifically, we design a Local and Global Feature Interaction Module (LGFI) to adaptively fuse global facial and local features with different receptive fields obtained by convolutions. In LGFI, self-attention is responsible for extracting global features, our proposed Residual Depth Feature Extraction Module (RDFE) extracts local features at different scales using separable convolutional kernels of different sizes, and our proposed Selective Kernel Attention Fusion Module (SKAF) is responsible for weighted fusion of these two parts of features for our model to adaptively perform selective fusion during training. In addition, we also utilize our proposed SKAF as a crucial fusion module in our Encoder and Decoder Feature Fusion Module (EDFF) to further perform feature communications of our method by fusing features at different scales from the encoder-decoder processes.
@@
In the following sections, we will introduce related work and discuss their disadvantages in Section II. Then, we will introduce our method and proposed core modules in Section III. Next, we will evaluate our method and analyze our ablation experiments in Section IV. Finally, we will conclude this paper according to our experiments in Section IV-E.
@@ II-A Face Super-Resolution
Attention-based FSR methods have been proposed to promote FSR to avoid the adverse effects of inaccurate prior estimates on FSR. Zhang et al. [16] proposed a supervised pixel-by-pixel generation of the adversarial network to improve face recognition performance during FSR. Chen et al. [17] proposed the SPARNet, which can focus on important facial structure features adaptively by using spatial attention in residual blocks. Lu et al. [18] proposed a partial attention mechanism to enhance the consistency of the fidelity of facial detail and facial structure. Bao et al. [19] introduced the equalization texture enhancement module to enhance the facial texture detail through histogram equalization. Wang et al. [20] critical introduced Fourier transform into FSR, fully exploring the correlation between spatial domain features and frequency domain features. Shi et al. [21] designed a two-branch network, which introduces convolution based on local changes to enhance the ability of convolution. Liu et al. [10] improves the interaction ability of regional and global features through designed hybrid attention modules. Li et al. [22] designed a wavelet-based network to reduce the loss of downsampling in the encoder-decoder. Although the above methods can reconstruct reasonable FSR images, they cannot promote the efficient fusion of local features with global features and different features at different stages of the encoder-decoder, affecting FSR’s efficiency and accuracy.
@@ III-A1 Encoding stage
The encoding stage in our network aims to extract facial features of different scales. In this stage, given a degraded face image , first, a convolution is used to extract shallow facial features. Then, extracted facial features are further refined by three encoder stages. Each encoder comprises our designed Local and Global Feature Interaction Module (LGFI) and a downsampling operator. After each encoder, the input face feature’s channel counts will be doubled, and the size of the image of the input face feature will be halved. As shown in Fig. 2, the features obtained after three encoders are as follows: , , .
@@ III-B2 Residual depth feature extraction module (RDFE)
III-B3 Selctive Kernel Attention Fusion Module (SKAF)
@@
where represents convolution, represents convolution, indicates the concat operation along the channel dimension. Then, we impose pooling to learn the weight of obtained hybrid features, where the weight reflects the importance of features under different receptive fields. The process of obtaining the weight for selecting required facial features is as follows:
where , and indicates the average and max pooling operation along the channel direction and sigmoid function, respectively. Finally, we multiply the weights obtained from the above calculations with the local and global features, respectively. Thus, our SKAF can adaptively select the important local and global information required for reconstruction. The process of obtaining important local and global features , by adaptive weight selection can be expressed as
@@ III-C Encoder and Decoder Feature Fusion Module (EDFF)
To fully utilize the multi-scale features extracted from the encoding and decoding stage, we introduce an EDFF to fuse different features, enabling our AMINet with better feature propagation and representation capabilities. As shown in Fig. 2, our EDFF mainly utilizes our proposed SKAF to fuse and select different scale features required for reconstruction. Given the feature the feature from the decoding stage and the encoding stage, respectively. Firstly, we concatenate features from the encoding and decoding stages along the channel dimension. Then, a convolution is used to reduce the channel counts and reduce the process’s computational costs to obtain two weights through our SKAF, which can be expressed as
where represents Selective Kernel Attention Fusion Module, stands for the convolutional layer, and denotes the operation of concatenating features across the channel dimension. Next, we feed the obtained two weights into two branches for multiplication. Through this operator, we obtain the selected facial features from hybrid features obtained by fusing the encoding and decoding features. Finally, we add the features of the two branches. The process is denoted as
@@ IV-C2 Comparison between LGFI and Transformer
As shown in Fig. 4, LGFI uses a dual branch structure to interact with representing the local and global features. In contrast, the traditional Transformer in Restormer [55] uses a serial structure to link the local and global features. To verify the effectiveness of LGFI, we replace all LGFIs in the network with Transformers and conduct comparative experiments with similar parameters between the two models. From Table II, we can see that the network’s performance using LGFI is better when the two networks maintain similar parameters. This is because LGFI utilizes the features of both local and global branches for interaction, facilitating the communication of multi-scale facial information.
@@ IV-C3 Comparison between RDFE and FFN
The feed-forward network (FFN) performs independent nonlinear transformations of the inputs at each position to help the Transformer capture local features, but it lacks the ability to extract multi-scale features, which is not favorable for accurate FSR. In contrast, our RDFE can extract multi-scale local features well. To compare RDFE and FFN, we replace RDFE with FFN while keeping the parameters of the two models similar. As shown in Table III, since FFN’s ability to capture feature interactions is limited compared to our RDFE that utilizes multiple branches to capture different receptive field facial features, our RDFE performs much better than FFN with similar computational consumption.
@@ IV-C5 Effectiveness of SKAF
SKAF is an important component of LGFI, facilitating information exchange between local and global branches. We perform ablation experiments to validate our SKAF module’s impact and assess the combined approach’s practicality. Since SKAF consists of dual branch convolutional layers, maximum pooling layers, and average pooling layers, we verify the effectiveness of module components in SKAF. From Table V, we have the following observations: (a) From the last three rows of the table, we find that using a single pooling branch results in reduced performance, while using average pooling alone results in lower performance than using maximum pooling alone. This is because the salient features of the face are the key to facial recovery, with maximum pooling focusing on salient facial feature information. In contrast, average pooling focuses on the overall information of the face. (b) Compared to the third and fifth rows, it can be concluded that using both and simultaneously can improve performance and fully utilize key facial information under different receptive fields.
@@ IV-D1 Comparison on CelebA dataset
We conduct a quantitative comparison of AMINet against existing FSR methods on the CelebA test set, as detailed in Table VII. Our AMINet outperforms all evaluation metrics, including PSNR, SSIM, LPIPS, and VIF, which fully demonstrates its efficiency. This strongly validates the effectiveness of AMINet. Additionally, the visual comparison in Fig. 7 reveals that previous FSR methods struggled to reproduce facial features like the eyes and mouth accurately. In contrast, AMINet excels at preserving the facial structure and producing more precise results, proving its effectiveness.
@@ IV-D2 Comparison on Helen dataset
We evaluate our method on the Helen test set to further assess AMINet’s versatility. Table VII provides a quantitative comparison of 8 FSR results about it, where AMINet achieves the better performance. Visual comparisons in Fig. 8 indicate that existing FSR methods struggle to maintain accuracy, leading to blurred shapes and a loss of facial details. In contrast, AMINet successfully preserves facial contours and details, reinforcing its effectiveness and adaptability across different datasets. In addition, we also provide a visual comparison of the face parsing for the recovered images, as shown in Fig. 9, which clearly shows that our AMINet can restore accurate components.
@@ IV-E Model Complexity Analysis
In addition to the performance indicators mentioned earlier, the number of model parameters, inference time, and computational complexity are crucial factors in evaluating performance. As shown in the Table X, we have selected some models for comparison. Meanwhile, as shown in Fig. 1, we compare our model with existing ones in terms of parameters, PSNR values, and inference speed. We can see that AMINet can have faster inference time, smaller parameter count, and computational complexity with similar performance to SCTANet. In addition, we provide a small parameter version called ”AMINet-S”. AMINet-S performs similarly to SFMNet in terms of parameter quantity, while AMINet has more advantages in computational complexity and inference time. This is thanks to AMINet achieving high efficiency by promoting extensive multi-scale feature exchange within the network. This design enables features from different receptive fields to effectively interact, allowing the model to adaptively select the most relevant information for facial reconstruction. Therefore, AMINet maintains strong facial super-resolution performance while balancing parameters and computational costs.

### v3

@@ Abstract
Recently, CNN and Transformer hybrid networks demonstrated excellent performance in face super-resolution (FSR) tasks. Since numerous features at different scales in hybrid networks, how to fuse these multi-scale features and promote their complementarity is crucial for enhancing FSR. However, existing hybrid network-based FSR methods ignore this, only simply combining the Transformer and CNN. To address this issue, we propose an attention-guided Multi-scale interaction network (AMINet), which contains local and global feature interactions and encoder-decoder phase feature interactions. Specifically, we propose a Local and Global Feature Interaction Module (LGFI) to promote fusions of global features and different receptive fields’ local features extracted by our Residual Depth Feature Extraction Module (RDFE). Additionally, we propose a Selective Kernel Attention Fusion Module (SKAF) to adaptively select fusions of different features within LGFI and encoder-decoder phases. Our above design allows the free flow of multi-scale features from within modules and between the encoder and decoder, which can promote the complementarity of different scale features to enhance FSR. Comprehensive experiments confirm that our method consistently performs well with less computational consumption and faster inference.
@@
However, while existing hybrid-based FSR methods consider utilizing features from different scales to improve FSR, they ignore the problem of how to fuse these multi-scale features to make their properties better complement each other. For example, Faceformer [8] simply parallelizes the connected CNN modules and the window-based Transformer [9] modules. SCTANet [10] also only juxtaposes spatial attention-based residual blocks and multi-head self-attention in designed modules. CTCNet [5] simply connects the CNN module in tandem with the Transformer module. These methods overlook the importance of blending multi-scale features in a complementary manner and enabling smooth information flow across different scales to refine facial details effectively.
To address this problem, we propose an Attention-Guided Multi-scale Interaction Network (AMINet) for FSR in this work. Our AMINet fuses multi-scale features in two main ways, including the fusion of features obtained from self-attention and convolution, and the fusion of features at different stages of the encoder-decoder. Specifically, we design a Local and Global Feature Interaction Module (LGFI) to adaptively fuse global facial and local features with different receptive fields obtained by convolutions. In LGFI, self-attention is responsible for extracting global features, our proposed Residual Depth Feature Extraction Module (RDFE) extracts local features at different scales using separable convolutional kernels of different sizes, and our proposed Selective Kernel Attention Fusion Module (SKAF) is responsible for weighted fusion of these two parts of features for our model to adaptively perform selective fusion during training. In addition, we also utilize our proposed SKAF as a crucial fusion module in our Encoder and Decoder Feature Fusion Module (EDFF) to further perform feature communications of our method by fusing features at different scales from the encoder-decoder processes.
@@
In the following sections, we will introduce related work and discuss their disadvantages in Section II. Then, we will introduce our method and proposed core modules in Section III. Next, we will evaluate our method and analyze our ablation experiments in Section IV. Finally, we will discuss our method in Section V and conclude this paper in Section VI.
@@ II-A Face Super-Resolution
Attention-based FSR methods have been proposed to promote FSR to avoid the adverse effects of inaccurate prior estimates on FSR. Zhang et al. [16] proposed a supervised pixel-by-pixel generation of the adversarial network to improve face recognition performance during FSR. Chen et al. [17] proposed the SPARNet, which can focus on important facial structure features adaptively by using spatial attention in residual blocks. Lu et al. [18] proposed a partial attention mechanism to enhance the consistency of the fidelity of facial detail and facial structure. Bao et al. [19] introduced the equalization texture enhancement module to enhance the facial texture detail through histogram equalization. Wang et al. [20] critical introduced Fourier transform into FSR, fully exploring the correlation between spatial domain features and frequency domain features. Shi et al. [21] designed a two-branch network, which introduces a convolution based on local changes to enhance the ability of the convolution. Liu et al. [10] improves the interaction ability of regional and global features through designed hybrid attention modules. Li et al. [22] designed a wavelet-based network to reduce the loss of downsampling in the encoder-decoder. Although the above methods can reconstruct reasonable FSR images, they cannot promote the efficient fusion of local features with global features and different features at different stages of the encoder-decoder, affecting FSR’s efficiency and accuracy.
@@ III-A1 Encoding stage
The encoding stage in our network aims to extract facial features of different scales. In this stage, given a degraded face image , first, a convolution is used to extract shallow facial features. Then, extracted facial features are further refined by three encoder stages. Each encoder comprises our designed Local and Global Feature Interaction Module (LGFI) and a downsampling operator. After each encoder, the input face feature’s channel count will be doubled, and the size of the image of the input face feature will be halved. As shown in Fig. 2, the features obtained after three encoders are as follows: , , .
@@ III-B2 Residual depth feature extraction module (RDFE)
III-B3 Selective Kernel Attention Fusion Module (SKAF)
@@
where represents convolution, represents convolution, indicates the concat operation along the channel dimension. Then, we impose pooling to learn the weight of the obtained hybrid features, where the weight reflects the importance of features under different receptive fields. The process of obtaining the weight for selecting required facial features is as follows:
where , and indicate the average and max pooling operation along the channel direction and sigmoid function, respectively. Finally, we multiply the weights obtained from the above calculations with the local and global features, respectively. Thus, our SKAF can adaptively select the important local and global information required for reconstruction. The process of obtaining important local and global features , by adaptive weight selection can be expressed as
@@ III-C Encoder and Decoder Feature Fusion Module (EDFF)
To fully utilize the multi-scale features extracted from the encoding and decoding stage, we introduce an EDFF to fuse different features, enabling our AMINet with better feature propagation and representation capabilities. As shown in Fig. 2, our EDFF mainly utilizes our proposed SKAF to fuse and select different scale features required for reconstruction. Given the feature , the feature from the decoding stage and the encoding stage, respectively. Firstly, we concatenate features from the encoding and decoding stages along the channel dimension. Then, a convolution is used to reduce the channel counts and reduce the process’s computational costs to obtain two weights through our SKAF, which can be expressed as
where represents Selective Kernel Attention Fusion Module, stands for the convolutional layer, and denotes the operation of concatenating features across the channel dimension. Next, we feed the two obtained weights into two branches for multiplication. Through this operator, we obtain the selected facial features from hybrid features obtained by fusing the encoding and decoding features. Finally, we add the features of the two branches. The process is denoted as
@@ IV-C2 Comparison between LGFI and Transformer
As shown in Fig. 4, LGFI uses a dual-branch structure to interact with representing the local and global features. In contrast, the traditional Transformer in Restormer [55] uses a serial structure to link the local and global features. To verify the effectiveness of LGFI, we replace all LGFIs in the network with Transformers and conduct comparative experiments with similar parameters between the two models. From Table II, we can see that the network’s performance using LGFI is better when the two networks maintain similar parameters. This is because LGFI utilizes the features of both local and global branches for interaction, facilitating the communication of multi-scale facial information.
@@ IV-C3 Comparison between RDFE and FFN
The feed-forward network (FFN) performs independent nonlinear transformations of the inputs at each position to help the Transformer capture local features, but it lacks the ability to extract multi-scale features, which is not favorable for accurate FSR. In contrast, our RDFE can extract multi-scale local features well. To compare RDFE and FFN, we replace RDFE with FFN while keeping the parameters of the two models similar. As shown in Table III, since FFN’s ability to capture feature interactions is limited compared to our RDFE which utilizes multiple branches to capture different receptive field facial features, our RDFE performs much better than FFN with similar computational consumption.
@@ IV-C5 Effectiveness of SKAF
SKAF is an important component of LGFI, facilitating information exchange between local and global branches. We perform ablation experiments to validate our SKAF module’s impact and assess the combined approach’s practicality. Since SKAF consists of dual-branch convolutional layers, maximum pooling layers, and average pooling layers, we verify the effectiveness of module components in SKAF. From Table V, we have the following observations: (a) From the last three rows of the table, we find that using a single pooling branch results in reduced performance, while using average pooling alone results in lower performance than using maximum pooling alone. This is because the salient features of the face are the key to facial recovery, with maximum pooling focusing on salient facial feature information. In contrast, average pooling focuses on the overall information of the face. (b) Compared to the third and fifth rows, it can be concluded that using both and simultaneously can improve performance and fully utilize key facial information under different receptive fields.
@@ IV-D1 Comparison on CelebA dataset
We conduct a quantitative comparison of AMINet against existing FSR methods on the CelebA test set, as detailed in Table VII. Our AMINet outperforms all evaluation metrics, including PSNR, SSIM, LPIPS, and VIF, which fully demonstrates its efficiency. This strongly validates the effectiveness of AMINet. Additionally, the visual comparison in Fig. 7 reveals that previous FSR methods struggled to reproduce facial features like the eyes and mouth accurately. In contrast, AMINet excels at preserving the facial structure and producing more precise results.
@@ IV-D2 Comparison on Helen dataset
We evaluate our method on the Helen test set to further assess AMINet’s versatility. Table VII provides a quantitative comparison of 8 FSR results about it, where AMINet achieves the better performance. Visual comparisons in Fig. 8 indicate that existing FSR methods struggle to maintain accuracy, leading to blurred shapes and a loss of facial details. In contrast, AMINet successfully preserves facial contours and details, reinforcing its effectiveness and adaptability across different datasets. In addition, we provide a visual comparison of face parsing maps for recovered face images, as shown in Fig. 9, which clearly shows that our AMINet facilitates downstream tasks such as face parsing maps segmentation.
@@ IV-E Model Complexity Analysis
In addition to the performance indicators mentioned earlier, the number of model parameters, inference time, and computational complexity are crucial factors in evaluating performance. As shown in the Table X, we have selected some models for comparison. Meanwhile, as shown in Fig. 1, we compare our model with existing ones in terms of parameters, PSNR values, and inference speed. We can see that AMINet can have faster inference time, smaller parameter count, and computational complexity with similar performance to SCTANet. In addition, we provide a small parameter version called ”AMINet-S”. AMINet-S performs similarly to SFMNet in terms of parameter quantity, while AMINet has more advantages in computational complexity and inference time. This is thanks to AMINet achieving high efficiency by promoting extensive multi-scale feature exchange within the network. This design enables features from different receptive fields to interact effectively, allowing the model to select the most relevant information for facial reconstruction adaptively. Therefore, AMINet maintains strong FSR performance while balancing parameters and computational costs.